# From Decision to Action in Surgical Autonomy: Multi-Modal Large Language Models for Robot-Assisted Blood Suction

### v1

@@ Abstract
Index Terms — Multi-modal LLMs, robot-assisted surgery, blood suction, decision-making, surgical autonomy.
@@ I Introduction
Autonomous execution of surgical sub-tasks, which are generally learned through model-based methods [13], or data-driven approaches such as deep reinforcement learning (DRL) [14] and imitation learning [15, 16] often leads to behavior that is not fully interpretable by humans and lacks explainability and adaptability in decision-making. These learning methods often prioritize maximizing cumulative rewards and the exploration of state space efficiently, and do not implicitly account for the risks of the actions taken; hence, in their current forms, they cannot assure any safety standards [17, 18, 19]. Also, DRL and imitation learning lack interpretability and struggle with dynamic adaptability. For instance, in scenarios where surgical conditions deviate from the norm due to unexpected patient anatomy or sudden complications such as bleeding, these systems might not adequately adjust their strategies.
Additionally, the ability to reason and take action in response to an unforeseen and unexpected circumstance cannot be easily accounted for solely in the reward function of a DRL-trained agent. This gap underscores the need for a reasoning framework that can adapt to unforeseen challenges, ensuring efficacy and transparent decision-making in autonomous surgical tasks. As surgical sub-tasks involve physical interaction with patients, any automation and decision-making by the robot must be clear and understandable by the operator and resemble human-like behavior to ensure the safety and reliability of the procedures.
@@
Autonomous surgical planning in robot-assisted surgical sub-tasks demands a human-like reasoning unit that would pre-operatively plan out the execution of the task and intraoperatively modify the plan to accommodate unforeseen circumstances. This would enhance the explainability of the decisions made by the robot and minimize the potential risks associated with them. Large Language Models (LLMs), trained on vast amounts of text data, have revolutionized the fields of natural language understanding and have been adopted in various research domains beyond natural language processing, such as planning and interaction for robots. They can serve as a high-level reasoning unit that would break down a given command into smaller subtasks for it to be passed on to the lower levels of the robot for subtask execution, including motion planning and control.
@@
Integrating LLMs into robots poses challenges due to their struggle with real-world complexities. While LLMs offer general knowledge and expertise, they lack a connection to tangible reality, leading to errors and potentially unsafe recommendations. Extracting information from LLMs for robots requires balancing theoretical knowledge with practical understanding to navigate dynamic environments and lead to effective human-robot interaction. An agent would need to comprehend semantic aspects of the world: the range of skills available, how these skills influence the world, and how changes to the world map back to the language [20]. This integration of robotics and LLMs, which have been introduced through Google’s SayCan [21], PaLM-E[22], and more recently explored methods [23, 24, 25, 26, 27, 28, 29, 30], presents an immense opportunity for exploration in various domains such as surgical robotics.
@@
By leveraging multi-modal LLMs, in this work we aim to surmount the limitations of current autonomous systems, introducing a level of reasoning and adaptability previously unattainable in robot-assisted surgeries. This integration is pioneering, as it combines the theoretical knowledge of LLMs with the practical demands of surgical environments as identified and communicated by the medical staff. We propose an LLM-powered framework capable of high-level reasoning, mid-level motion planning, and execution for autonomous blood suctioning in robot-assisted surgeries. The reasoning and prioritization responsibility is delegated to the higher-level task planning LLM, and the motion planning and execution are delegated to the DRL model, leading to a distributed agency between the two components. Expanding on the foundation laid with autonomous blood suctioning, our vision extends beyond this surgical subtask and sets a precedent for the future of autonomous systems in the surgical field.
The main contribution of this work is as follows: 1) We propose an LLM-powered framework for autonomous robot-assisted blood suctioning, 2) We compare the performance of LLM reasoning to random reasoning and no reasoning modules in blood removal time and tool movement, 3) We analyze how augmenting the prompts with context and expert-defined guidelines changes the reasoning capabilities of the LLM in zero-shot prompting. A user study is also performed to assess the closeness to human-like behavior among the three modules.
The paper is structured as follows. Section \@slowromancapii@ reviews the integration of LLMs in robotics and surgery and how it has influenced the two. Section \@slowromancapiii@ introduces the methods used, including the architecture of the system, multi-modal LLMs and prompt augmentation, the simulation environment, and the deep reinforcement learning module. In Section \@slowromancapiv@, the experiment results are presented and discussed. Section \@slowromancapv@ concludes the paper and explains the limitations and future work.
@@ II-B Large Language Models in Surgery and Surgical Planning
Although an initial step towards integrating LLMs in robotic surgery has been made in SUIFA [35] though tasks such as needle lift and shunt insertion, the current gaps in the literature include the adoption of LLMs as reasoning units in the planning process of surgical tasks where decision-making is crucial. In this work, we investigate the reasoning capability of LLMs in prioritizing the suction of multiple blood pools under different circumstances and integrate them into robot-assisted surgery for autonomous blood suctioning.
@@ III-A Blood Suctioning Task and System Architecture
The proposed system architecture can be seen in Fig. 2. An image of the tissue scene containing multiple blood pools is contoured and passed on to the multi-modal large language model, along with a text prompt asking the LLM to prioritize and suction the blood pools. The multi-modal LLM, as the high-level reasoning unit, uses its reasoning capability to prioritize the order in which the blood pools need to be suctioned and accordingly informs the masking sensor. The pools are masked in the order that they need to be suctioned and fed into the trained DRL agent, which acts as the lower-level motion planner, along with the tissue height map, and leads to an action taken by the suction tool. In complex scenarios where additional information is needed for the LLM to prioritize the blood suction, the human can augment the initial prompt and provide context. An image of the input prompt, response by the LLM, and timesteps of executing is shown in Fig. 1.
@@ III-B Multi-Modal Large Language Models
Our methodology leverages multi-modal LLMs to allow for information from diverse modalities, including text and images. In multi-modal LLMs, textual data undergoes processing through a standard language model architecture. Feature embeddings extracted from image inputs are concatenated with the textual embeddings, yielding a multi-input representation. This fused representation is then fed into a multi-layer neural network, facilitating joint learning across modalities. Multi-modal LLMs serve as robust tools for image understanding and reasoning. In this work, we employed the pre-trained OpenAI GPT-4V model [36] for image understanding and reasoning.
As shown in Fig. 1 and explained in the system architecture section, the multi-modal LLM accepts a text prompt and an image of the contoured and labeled blood pools. The prompt outlines essential details about blood pools, including the presence of clots, and signs of active bleeding. Each of these aspects is crucial in deciding the urgency of suctioning blood pools during medical procedures. Size indicates the amount of blood, while clots indicate coagulation and cessation of blood flow. Active bleeding is a significant signal that could suggest the potential for blood to spread more extensively. By incorporating these factors into the prompt, we equip the multi-modal LLM with essential context for making informed decisions regarding blood pool suctioning. The LLM conducts reasoning based on the image and the prompt, yielding the proper priority for suctioning the blood, and explains why this priority is chosen in a zero-shot manner.
Zero-shot and few-shot approaches extend the model’s abilities to comprehend and generate text beyond the training data. Zero-shot learning in LLMs involves the model generating text for prompts it has not been explicitly trained on. This is achieved by leveraging its understanding of language semantics and structures acquired during training, thereby producing relevant responses even for unseen prompts. Conversely, few-shot learning entails fine-tuning the LLM with a limited number of examples per prompt. By providing a small amount of labeled data, the model can adapt its parameters to better understand and generate text specific to the given prompt, resulting in more contextually relevant responses.
@@
III-C Blood Suction Environment and Mask Sensor Mechanism
In our recent work [10], a blood suction simulation environment for RL was built using position-based fluids (PBF) based on Nvidia PhysX 5, Unity, and Unity’s ML-Agents toolkit. PBF is an approach that represents fluids using a large number of small particles that interact with each other. PhysX is a real-time physics engine with GPU optimization, which allows for PBF simulation. With PhysX 5 as the low-level physics engine, the main simulation environment was built in Unity, a 3D game development software. We leverage the existing simulation environment and further build upon it in this work.
The simulation environment consists of a randomly generated tissue that contains the blood, the simulated blood, and a suction tool. We simulated a fixed amount of blood (4000 particles) through PBD and added suction force to each particle within a suction range to simulate the effect of suction. Particles that are suctioned close enough to the suction tool will be removed and marked as inactive. To introduce randomness in the shape of the tissue, Bezier surfaces with random control points were used to generate random shapes. The Bezier surfaces are represented by
[…]
$\ B_{n,i}(x)=\frac{n!}{i!\cdot(n-i)!}\cdot x^{i}\cdot(1-x)^{n-i}.$ (2)
In this work, several features are added to the simulation environment, including a module that would allow blood to continuously add to a random pool representing active bleeding, the addition of a capsule-shaped object representing a blood clot that would be randomly positioned on the tissue, and the ability to generate multiple independent blood pools. We also developed a blood pool detector algorithm that takes in the raw image of the scene along with the suction orders from the multi-modal LLM and outputs the mask of the target blood pool based on the priority reasoned by the LLM. This mechanism would allow the agent to plan the motion for suction of only the target pool and would blind the agent to other existing pools unless commanded otherwise by the LLM.
We developed four simulation environments. In Environment 1, as can be seen in Fig. 3, four independent pools are randomly generated without the presence of a blood clot or active bleeding. Environment 2 features the introduction of a capsule-shaped object representing a blood clot, randomly generated near one of the pools, indicating coagulation and the cessation of potential bleeding. In Environment 3, active bleeding occurs randomly in one of the pools for a fixed time interval, with no sign of a blood clot. Finally, Environment 4 includes a pool actively bleeding, while another contains a blood clot. To test the realism of our simulation environments, visual comparisons with real surgical environments were conducted, ensuring our model visually reflects the dynamics of blood flow and bleeding during real surgical procedures.
@@ III-D Motion Planning Using Deep Reinforcement Learning
In the above equation, $t$ is the time step, $N_{p}^{t}$ is the number of active particles, $\delta(N_{p}^{t+1})$ denotes whether there are active particles remaining, and $\|a_{t}\|$ is the norm of the actions. $C_{1}=5$ and $C_{2}=0.02$ are two weighting factors. The observation includes the tissue height map, the binary image mask of the blood (stacked with 3 from previous steps), and the suction tool location (stacked with 4 from previous steps). The binary image mask of the blood is synthesized from the current positions of all active particles in the blood.
@@ IV Experiments and Results
We investigate four reasoning modules in our experiments. When the LLM, as the high-level reasoning unit, reasons the sequence for suctioning blood pools without any additional context, it is termed LLM Reasoning Without Context (LRWOC). If additional context is provided by the assistant, leading to an augmented prompt for the LLM, we call this LLM Reasoning With Context (LRWC). When the DRL agent receives a randomly generated order based on a random permutation of the number of blood pools, this is known as Random Reasoning (RR). If the DRL agent proceeds to suction blood pools solely based on its reward function without input from a higher-level unit, this is termed No Reasoning (NR).
@@ IV-A Comparison of LLM Reasoning with Random Reasoning and No Reasoning
To evaluate the performance of the RR, NR, LRWOC, and LRWC modules, we simulated the blood suction task across 400 distinct scenes (100 scenes per environment). Fig. 5 illustrates the blood remaining percentage over time, providing a comparative analysis of the different reasoning modules across different environments. Additionally, Table I presents key metrics such as the mean and standard deviations of the time to suction 50% of the blood ($T_{50}$), the area under the curve (AUC), and suction tool movement.
@@
In Environment 2 (Fig. 5b), where active bleeding is present in one of the pools, the LLM reasoning module gives priority to suctioning the pool with active bleeding first, even if it is smaller than others. This approach results in a slower start when compared to NR and RR modules. After addressing the pool with active bleeding, LR then proceeds to prioritize suctioning the pools based on their size, similar to the strategy observed in Environment 1. Providing additional instructions as the augmented prompt did not change the LLM reasoning in Environments 1 and 2 and the focus remained on prioritizing active bleeding, followed by pool size. This resulted in the same LRWOC and LRWC, denoted as LR.
In Environment 3 (Fig. 5c), the presence of a blood clot adds complexity to suctioning. LRWOC tends to prioritize the suction of the pool with the blood clot potentially due to its perceived complexity. However, an expert can define rules for the LLM to follow, tailoring its reasoning to specific conditions. For this environment, we establish the following rule: 1) Address active bleeding first, 2) Consider pool size next, and 3) Address the blood clot pool last, as coagulation ensures that flow in this pool has ceased and will not propagate further. An example of this is shown in Fig. 4. This context leads to a faster average $T_{50}$ in LRWC, prioritizing blood pool size after active bleeding. In Environment 4 (Fig. 5d), both LRWOC and LRWC modules start by suctioning actively bleeding pools. LRWC prioritizes larger pools next and saves the pool with the blood clot for last, resulting in a faster $T_{50}$ compared to LRWOC, which targets the pool with the blood clot before proceeding based on pool size.
TABLE I: Key metrics in comparison of reasoning modules expressed as mean±$\pm$±std. (Time step = 0.02 seconds)
$T_{50}$ AUC Tool Movement
Environment 1 RR 225$\pm$82 26467$\pm$5336 33.9$\pm$3.2
NR 164$\pm$38 20547$\pm$7086 32.4$\pm$4.1
LR 145$\pm$30 21016$\pm$3814 32.5$\pm$2.8
Environment 2 RR 223$\pm$89 25608$\pm$5641 33.3$\pm$3.4
NR 154$\pm$37 19452$\pm$5530 31.5$\pm$3.9
LR 161$\pm$28 20998$\pm$3246 33$\pm$2.8
Environment 3 RR 222$\pm$71 25958$\pm$4146 32.4$\pm$2.9
NR 166$\pm$34 19303$\pm$4410 31.2$\pm$4.2
LRWOC 239$\pm$58 24887$\pm$3268 32.4$\pm$3.0
LRWC 151$\pm$23 22540$\pm$3147 32.7$\pm$2.9
Environment 4 RR 227$\pm$86 25612$\pm$4707 33.7$\pm$3.2
NR 155$\pm$34 18843$\pm$5116 30.4$\pm$3.8
LRWOC 179$\pm$55 21885$\pm$4114 33.7$\pm$2.9
LRWC 137$\pm$18 19629$\pm$3294 32.9$\pm$2.8
@@
Figure 6: Time to suction 50% ($T_{50}$) of the blood in different environments.
Fig. 6 illustrates the time to suction 50% of the blood ($T_{50}$) and shows statistically significant improvement (*) in almost all environments when LLM reasoning is used and context is provided. However, in the surgical context, speed and tool movement should not always be prioritized, especially in the presence of critical complexities such as active bleeding. The LLM’s ability to consider and prioritize addressing these complexities is valuable for the advancement of decision-making in autonomous surgical robots.
@@
Table I shows that the LRWC module results in smaller standard deviations across all metrics and environments, indicating more consistent performance. Although the NR module exhibited a marginally smaller average suction tool movement and AUC, its inability to reason and adapt to unforeseen circumstances makes it unreliable in highly dynamic surgical settings.
IV-B User Study on Closeness to Human-Like Behaviour

### v2

@@ Abstract
Index Terms:
Medical robots and systems, multi-modal large language models, surgical robot, planning, laparoscopy.
[…]
@@ I Introduction
Autonomous execution of surgical sub-tasks is typically learned through model-based methods [13] or data-driven approaches like deep reinforcement learning (DRL) [14] and imitation learning [15, 16]. However, these methods often lead to behavior that lacks human interpretability, explainability, and adaptability in decision-making. They prioritize maximizing cumulative rewards and efficient exploration of the state space but do not implicitly account for the risks of the actions taken, thus failing to assure safety standards [17, 18, 19]. Moreover, DRL and imitation learning struggle with dynamic adaptability, particularly in scenarios where surgical conditions deviate from the norm due to unexpected patient anatomy or sudden complications such as bleeding. This inability to reason and act in unforeseen circumstances underscores the need for a reasoning framework that can adapt to challenges, ensuring efficacy and transparent decision-making in autonomous surgical tasks. As surgical sub-tasks involve physical interaction with patients, any automation and decision-making by the robot must be clear and understandable by the operator, resembling human-like behavior to ensure safety and reliability.
@@
Autonomous surgical planning in robot-assisted sub-tasks demands a human-like reasoning unit capable of pre-operatively planning task execution and intraoperatively modifying the plan to accommodate unforeseen circumstances. This approach would enhance the explainability of the robot’s decisions and minimize associated risks. Large Language Models (LLMs), trained on vast amounts of text data, have revolutionized natural language understanding and have been adopted in various domains beyond NLP, such as planning and interaction for robots. They can serve as a high-level reasoning unit, breaking down given commands into smaller subtasks to be executed by the robot’s lower-level systems, including motion planning and control.
@@
Integrating LLMs into robots poses challenges due to their struggle with real-world complexities. While LLMs offer general knowledge and expertise, they lack a connection to tangible reality, leading to errors and potentially unsafe recommendations. Extracting information from LLMs for robots requires balancing theoretical knowledge with practical understanding to navigate dynamic environments and facilitate effective human-robot interaction. An agent needs to comprehend semantic aspects of the world, the range of available skills, how these skills influence the environment, and how changes in the world map back to language [20]. The integration of robotics and LLMs, as introduced through Google’s SayCan [21], PaLM-E [22], and more recent methods [23, 24, 25, 26, 27, 28, 29, 30], presents immense opportunities for exploration in domains such as surgical robotics.
By leveraging multi-modal LLMs in a zero-shot manner, in this work we aim to surmount the limitations of current autonomous systems, introducing a level of reasoning and adaptability previously unattainable in robot-assisted surgeries. Integrating images with text allows the multi-modal LLM to capture important visual nuances, such as spatial relationships and the presence of surgical tools near blood pools, which may not be fully conveyed through text alone. This integration is pioneering, as it combines the theoretical knowledge of LLMs with the practical demands of surgical environments as identified and communicated by the medical staff. We propose an LLM-powered framework capable of high-level reasoning, mid-level motion planning, and execution for autonomous blood suctioning in robot-assisted surgeries. The reasoning and prioritization responsibility is delegated to the higher-level task planning LLM, and the motion planning and execution are delegated to the DRL model, leading to a distributed agency between the two components. Expanding on the foundation laid with autonomous blood suctioning, our vision extends beyond this surgical subtask and sets a precedent for the future of autonomous systems in the surgical field.
@@
The main contributions of this work are as follows:
  1. 1.
    We propose an LLM-powered framework for autonomous robot-assisted blood suction, where task reasoning and planning are managed by the LLM, while motion planning and execution are handled by a DRL-trained agent.
  2. 2.
    We compare the performance of LLM reasoning to random reasoning and no reasoning modules in terms of blood removal time and tool movement.
  3. 3.
    We analyze how augmenting the prompts with context and expert-defined guidelines influences the reasoning capabilities of the LLM in zero-shot prompting. A user study is also conducted to assess the similarity to human-like behavior across the three modules.
The paper is structured as follows. Section \@slowromancapii@ reviews the integration of LLMs in robotics and surgery and their impact on these fields. Section \@slowromancapiii@ introduces the methods used, including the system architecture, multi-modal LLMs, prompt augmentation, the simulation environment, and the deep reinforcement learning module. Section \@slowromancapiv@ presents and discusses the experiments and results. Section \@slowromancapv@ outlines the limitations and future work, while Section \@slowromancapvi@ concludes the paper.
@@ II-B Large Language Models in Surgery and Surgical Planning
Although an initial step towards integrating LLMs in robotic surgery has been made in SUIFA [35] through tasks such as needle lift and shunt insertion, the current gaps in the literature include the adoption of LLMs as reasoning units in the planning process of surgical tasks where decision-making is crucial. In this work, we investigate the reasoning capability of LLMs in prioritizing the suction of multiple blood pools under different circumstances and integrate them into robot-assisted surgery for autonomous blood suctioning.
@@ III-A Blood Suctioning Task and System Architecture
As shown in Fig. 1, an image of the tissue scene containing multiple blood pools is annotated with bounding boxes around each pool and passed to the multi-modal large language model, along with the text prompt, “Look at this image of multiple blood pools. Please prioritize and suction the blood effectively.” The proposed system architecture, illustrated in Fig. 2, depicts how the multi-modal LLM, as the high-level reasoning unit, uses its reasoning capability to prioritize the order in which the blood pools need to be suctioned and informs the masking sensor accordingly. The pools are masked in the order that they need to be suctioned and then fed into the trained DRL agent, which acts as the lower-level motion planner, along with the tissue height map, and leads to an action taken by the suction tool. In complex scenarios where additional information is needed for the LLM to prioritize suction, human input can augment the initial prompt with context.
@@ III-B Multi-Modal Large Language Models
Our methodology leverages multi-modal LLMs to allow for information from diverse modalities, including text and images. In multi-modal LLMs, textual data undergoes processing through a standard language model architecture. Feature embeddings extracted from image inputs are concatenated with the textual embeddings, yielding a multi-input representation. This fused representation is then fed into a multi-layer neural network, facilitating joint learning across modalities. In this work, we employed the pre-trained OpenAI GPT-4V model [36] for image understanding and reasoning.
As shown in Fig. 1 and explained in the system architecture section, the multi-modal LLM accepts a text prompt and an image of the labeled blood pools. The prompt outlines essential details about blood pools, including the presence of clots, and signs of active bleeding. Each of these aspects is crucial in deciding the urgency of suctioning blood pools during medical procedures. Size indicates the amount of blood, while clots indicate coagulation and cessation of blood flow. Active bleeding is a significant signal that could suggest the potential for blood to spread more extensively. By incorporating these factors into the prompt, we equip the multi-modal LLM with essential context for making informed decisions regarding blood pool suctioning. The LLM conducts reasoning based on the image and the prompt, yielding the proper priority for suctioning the blood, and explains why this priority is chosen in a zero-shot manner. Our zero-shot approach allows the LLM to generate relevant responses for each prompt without specific training examples, relying on its general understanding of language semantics and prompt cues to address previously unseen conditions effectively.
III-C Blood Suction Environment and Mask Sensor Mechanism
In our recent work [10], a blood suction simulation environment for RL was built using position-based fluids (PBF) based on Nvidia PhysX 5, Unity, and Unity’s ML-Agents toolkit. PBF is an approach that represents fluids using a large number of small particles that interact with each other. PhysX is a real-time physics engine with GPU optimization, which allows for PBF simulation. With PhysX 5 as the low-level physics engine, the main simulation environment was built in Unity. In this model, blood is simulated as particles influenced by forces like gravity and suction, with a spherical cone-shaped force field applied to particles near the suction tool. The force decays with distance to realistically simulate suction, removing particles once they reach a specified height threshold. We leverage the existing simulation environment and further build upon it in this work.
The simulation environment consists of a randomly generated tissue that contains the blood, the simulated blood, and a suction tool. We simulated a fixed amount of blood (4000 particles) through PBF and added suction force to each particle within a suction range to simulate the effect of suction. Particles that are suctioned close enough to the suction tool will be removed and marked as inactive. To introduce randomness in the shape of the tissue, Bezier surfaces with random control points were used to generate random shapes. The Bezier surfaces are represented by
[…]
$\ B_{n,i}(x)=\frac{n!}{i!\cdot(n-i)!}\cdot x^{i}\cdot(1-x)^{n-i}.$ (2)
@@ III-C Blood Suction Environment and Mask Sensor Mechanism
In this work, several features are added to the simulation environment, including a module that would allow blood to continuously add to a random pool representing active bleeding, the addition of a capsule-shaped object representing a blood clot that would be randomly positioned on the tissue, and the ability to generate multiple independent blood pools. We also developed a blood pool detector algorithm that takes in the raw image of the scene along with the suction orders from the multi-modal LLM and outputs the mask of the target blood pool based on the priority reasoned by the LLM. This mechanism would allow the agent to plan the suction motion of only the target pool and would blind the agent to other existing pools unless commanded otherwise by the LLM.
We developed four simulation environments. In Environment 1, as can be seen in Fig. 4, four independent pools are randomly generated without the presence of a blood clot or active bleeding. In Environment 2, active bleeding occurs randomly in one of the pools for a fixed time interval, with no sign of a blood clot. Environment 3 features the introduction of a capsule-shaped object representing a blood clot, randomly generated near one of the pools, indicating coagulation and the cessation of potential bleeding. Finally, Environment 4 includes a pool actively bleeding, while another contains a blood clot. To test the realism of our simulation environments, visual comparisons with real surgical environments were conducted, ensuring our model visually reflects the dynamics of blood flow and bleeding during real surgical procedures.
@@ III-D Motion Planning Using Deep Reinforcement Learning
In the above equation, $t$ is the time step, $N_{p}^{t}$ is the number of active particles, $\delta(N_{p}^{t+1})$ denotes whether there are active particles remaining, and $\|a_{t}\|$ is the norm of the actions. The weighting factors $C_{1}=5$ and $C_{2}=0.02$ were chosen to balance task efficiency and control stability, emphasizing blood removal while discouraging excessive motion. Specifically, $C_{1}$ is set to a higher value to prioritize the reward for full blood clearance, encouraging the agent to complete the task efficiently, while $C_{2}$ is relatively small, penalizing movements without restricting necessary adjustments. The observation includes the tissue height map, the binary image mask of the blood (stacked with 3 from previous steps), and the suction tool location (stacked with 4 from previous steps). The binary image mask of the blood is synthesized from the current positions of all active particles in the blood.
@@ IV Experiments and Results
We investigate four reasoning modules in our experiments. When the LLM, as the high-level reasoning unit, reasons the sequence for suctioning blood pools without any additional context, it is termed LLM Reasoning Without Context (LRWOC). If additional context is provided by the assistant, leading to an augmented prompt for the LLM, we call this LLM Reasoning With Context (LRWC). When the DRL agent receives a randomly generated order based on a random permutation of the number of blood pools, this is known as Random Reasoning (RR). If the DRL agent proceeds to suction blood pools solely based on its reward function without input from a higher-level unit, it is termed No Reasoning (NR).
@@ IV-A Comparison of LLM Reasoning with Random Reasoning and No Reasoning
To evaluate the performance of the RR, NR, LRWOC, and LRWC modules, we simulated the blood suction task across 400 distinct scenes (100 scenes per environment). Fig. 5 illustrates the blood remaining percentage over time, providing a comparative analysis of the different reasoning modules across different environments. Additionally, Table I presents key metrics such as the mean and standard deviations of the time to suction the actively bleeding pool in Environments 2 and 4 ($T_{AB}$), the time to suction 50% ($T_{50}$) and 95% ($T_{95}$) of the blood, and total tool path length (TTPL).
@@
Figure 6: Time to suction the actively bleeding pool ($T_{AB}$) in Environments 2 and 4 by different reasoning modules.
In Environment 2 (Fig. 5b), where active bleeding is present in one of the pools, the LLM reasoning module gives priority to suctioning the pool with active bleeding first, even if it is smaller than others. This approach results in a faster $T_{AB}$. After addressing the pool with active bleeding, the LLM then proceeds to prioritize suctioning the pools based on their size, similar to the strategy observed in Environment 1. Providing additional instructions as the augmented prompt did not change the LLM reasoning in Environments 1 and 2 and the focus remained on prioritizing active bleeding, followed by pool size. This resulted in the same LRWOC and LRWC, denoted as LR.
In Environment 3 (Fig. 5c), the presence of a blood clot adds complexity to suctioning. LRWOC tends to prioritize the suction of the pool with the blood clot potentially due to its perceived complexity. However, the user can define rules for the LLM to follow, tailoring its reasoning to specific conditions. For this environment, we establish the following rule: 1) Address active bleeding first, 2) Consider pool size next, and 3) Address the blood clot pool last, as coagulation ensures that flow in this pool has ceased and will not propagate further. An example of this is shown in Fig. 3. This context leads to a faster average $T_{50}$ in LRWC, prioritizing blood pool size after active bleeding. In Environment 4 (Fig. 5d), both LRWOC and LRWC modules start by suctioning actively bleeding pools. LRWC prioritizes larger pools next and saves the pool with the blood clot for last, resulting in a faster $T_{50}$ compared to LRWOC, which targets the pool with the blood clot before proceeding based on pool size.
TABLE I: Key metrics in comparison of reasoning modules expressed as mean±$\pm$±std (Time step = 0.02 seconds).
$T_{AB}$ $T_{50}$ $T_{95}$ TTPL
Environment 1 RR 225$\pm$82 570$\pm$128 33.9$\pm$3.2
NR 164$\pm$38 557$\pm$110 32.4$\pm$4.1
LR 145$\pm$30 554$\pm$101 32.5$\pm$2.8
Environment 2 RR 363$\pm$176 223$\pm$89 573$\pm$130 33.3$\pm$3.4
NR 390$\pm$174 154$\pm$37 514$\pm$106 31.5$\pm$3.9
LR 128$\pm$50 161$\pm$28 523$\pm$95 33$\pm$2.8
Environment 3 RR 222$\pm$71 563$\pm$124 32.4$\pm$2.9
NR 166$\pm$34 466$\pm$98 31.2$\pm$4.2
LRWOC 239$\pm$58 541$\pm$103 32.4$\pm$3.0
LRWC 151$\pm$23 553$\pm$100 32.7$\pm$2.9
Environment 4 RR 359$\pm$178 227$\pm$86 535$\pm$120 33.7$\pm$3.2
NR 393$\pm$172 155$\pm$34 462$\pm$104 30.4$\pm$3.8
LRWOC 130$\pm$49 179$\pm$55 509$\pm$93 33.7$\pm$2.9
LRWC 130$\pm$49 137$\pm$18 470$\pm$90 32.9$\pm$2.8
@@
Fig. 6 illustrates the time to suction the actively bleeding pool ($T_{AB}$) in Environments 2 and 4 and shows statistically significant improvement (*) when LLM reasoning is used. Table I shows that the LRWC module results in smaller standard deviations across all metrics and environments, indicating more consistent performance. Although the NR module exhibited a marginally smaller average TTPL, its inability to reason and adapt to unforeseen circumstances makes it unreliable in highly dynamic surgical settings.
IV-B Advantage of Multi-Modal LLMs
To further demonstrate the advantage of the multi-modal LLM, we conducted an experiment to test its ability to capture contextual details that may not be explicitly provided in text. In this experiment, we presented the model with 10 images in which a surgical tool was positioned near one of the blood pools, using our original prompt without any mention of the tool in text. The multi-modal LLM correctly recognized the presence of the tool in 80% of cases (8 out of 10) and incorporated this visual information into its decision-making process, placing this priority right after the active bleeding pool. Unlike hard-coded logic, which lacks flexibility in unforeseen scenarios, the LLM leverages context awareness and visual information to adapt to these conditions interpreting nuanced visual cues that a text-only input or a hard-coded logic might overlook.
IV-C User Study on Closeness to Human-Like Behaviour